What is the equation of a line with a slope of 4 and passing through the point (1,-1)? Thanks in advance!
Answer by ankor@dixie-net.com(22740)

What is the equation of a line with a slope of 4 and passing through the point (1,-1)?
:
Use the point/slope form: y - y1 = m(x - x1)
where
x1=1
y1=-1
m=4
:
y - (-1) = 4(x - 1)
y + 1 = 4x - 4
y = 4x - 4 - 1
y = 4x - 5
:
confirm this, substitute 1 for x and see that it will = -1